OpenStudy (anonymous):

Consider a spherical cow that is consuming a great deal of hay. He is fattening himself up at a rate of 200 liters/day, so how quickly is her radius growing (in cm/da) when he can't fit through the circular stall door (that's 2 meters wide)?

OpenStudy (ipwnbunnies):

I can help you with the first part. \[V(sphere) = (4/3)\pi r^{3}\]

OpenStudy (ipwnbunnies):

We're given dV/dt = 200 L/day; This problem is basically saying at what rate is the radius, r, increasing.

OpenStudy (anonymous):

Wait... 200 = 4(pI)(r)^2(dv/dt) right?

OpenStudy (ipwnbunnies):

I think (dr/dt) will be on the right side instead, because we're deriving the function with respect to t. Man, it's a little hazy to me. I hope someone else comes around to help out.

OpenStudy (ipwnbunnies):

(dV/dt) = 4pi*r^2*(dr/dt)

OpenStudy (anonymous):

200/40,000(pi) = 4(pi)(100)^2(dr/dt)?

OpenStudy (ipwnbunnies):

Ohhh, I know what I'm doing now. Let's forget about the derivative right now. We do need both parts of the problem to figure this out.

OpenStudy (anonymous):

Okay.

OpenStudy (ipwnbunnies):

So, the circular door is 2m wide. Therefore, when the cow's radius is over 1m, she can't fit through the door! We'll have to find the volume of the cow at 1m.

OpenStudy (ipwnbunnies):

We do need to do some conversions in the units though. The problem gives dV/dt in L/day, and they want dr/dt in cm/day. It's alright though. 1m * (100cm/1m) = 100 cm is the maximum radius of the cow. When we find the volume of the cow at 100 cm, units will be in cm^3.

OpenStudy (ipwnbunnies):

For dV/dt = 200 L/day = (200 L/day)(1000 mL/L)(1 cm^3 / mL)

OpenStudy (ipwnbunnies):

So, dV/dt = 200,000 cm^3/day

OpenStudy (ipwnbunnies):

NOW, we can find dr/dt using our derivative. dV/dt = 4*pi*r^2*(dr/dt)

OpenStudy (ipwnbunnies):

Solve for dr/dt, plug in the numbers and boom. The units will be in cm/day

OpenStudy (anonymous):

Is it 1/200(pi)?

OpenStudy (ipwnbunnies):

Sounds like I'm ranting, it takes alot of words to write out related rates problems. No, I got 5/pi

OpenStudy (ipwnbunnies):

Use dV/dt value that has units cm^3/day & use the radius value that is in cm.

OpenStudy (anonymous):

Remember it's 0.02 m right?

OpenStudy (ipwnbunnies):

No, that's the diameter.

OpenStudy (anonymous):

Okay, so it should be 5/pi?

OpenStudy (ipwnbunnies):

Should be...

OpenStudy (anonymous):

Okay, thanks.